What is the new volume, in liters, of a 25.0 L sample of air at 843.6...

  1. What is the new volume, in liters, of a 25.0 L sample of air at 843.6 mmHg and 152°C that is compressed and cooled to 35°C and 1520 mmHg?

2.. A balloon containing 7.83 moles of He has a volume of 27.5 L at a certain temperature and pressure. What would be the volume of the balloon, in liters, if 1.00 mole of Heis removed from the balloon at constant temperature and pressure?

3. Calculate the pressure in atmospheres, of 2.5 moles of F2 gas at 50°C, occupying a volume of 40.0 L

4. A sample of sulfur hexafluoride gas occupies 9.10 L at 198°C. Assuming that the pressure remains constant, what temperature (in °C) is needed to reduce the volume to 2.50 L?

5. An unknown gas with a mass of 4.52 g was found to occupy a volume of 875 mL at 1.69 atm and 65 °C. What is the molar mass of this gas

6. A gaseous mixture contains 25.0 g of each of the gases CO, CO2, and H2S. Assuming that the gases do not react with each other, calculate the mole fraction of each gas present in the mixture.
Answer #1

1)

According to ideal gas law.

PV=nRT where P is pressure ,V is volume ,n is no of moles,R is universal gas constant and T is temperature in Kelvin.

As n and R are constant.

So

(PV/T) = constant

This implies.(P1V1/T1)=(P2V2/T2)

P1= 843.6 mm-Hg V1= 25L and T1= (152+273)k = 425K.

P2= 1520 mm-Hg and T2 = (35+273)k = 308k

So ((843.6)(25)/(425)) = ((1520)(V2)/(308))

V2 comes out to be 10.06 L.

Answer is 10.06 litres